[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Nov 23, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Nov 24 02:50:20 UTC 2016


HSU Shift Log for Wed Nov 23 2016 21:50:14 GMT-0500 (EST)

Shift: R

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: Guidance submission for Otto timed out on the first try. It  
worked on the second try.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161124/a9dd1c69/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list